OpenStudy (anonymous):

Fan+Medal? Solve for h. P = mgh A. h = P – mg B.h=pg/m C.h=p/mg D. h = P + mg

OpenStudy (anonymous):

@I<3U

OpenStudy (anonymous):

@confluxepic

OpenStudy (kendricklamar2014):

C is the answer @Korey_Minter

OpenStudy (anonymous):

Thanks man!! COUld you help out on more @KendrickLamar2014

OpenStudy (anonymous):

Which literal equations are equivalent to b=a/h ? Choose all answers that are correct. A. h=bA B.h=a/b C.a=bh D.a=b/h @KendrickLamar2014

OpenStudy (confluxepic):

@Korey_Minter It would be better if you ask it in a different question. You can do that and tag me.

OpenStudy (confluxepic):

@Korey_Minter @~ What what the equation be if you multiply by h on both sides.

OpenStudy (confluxepic):

It would become (b)(h)=a.

OpenStudy (confluxepic):

Which is the same as a=bh.

OpenStudy (confluxepic):

And if you divide by b you will get h=a/b.

OpenStudy (confluxepic):

Those are the only correct answers.

OpenStudy (confluxepic):

@TheSmartOne @confluxepic

OpenStudy (confluxepic):

So B and C.